You are doing two projects P and Q. The probability that project P is
successful is equal to 2/3 and the probability that project Q is successful is equal to 4/5.
Whether or not these two projects are successful are independent of each other.
What is the probability that both P and Q are not successful?

Answer by stanbon(75887)

You are doing two projects P and Q. The probability that project P is
successful is equal to 2/3 and the probability that project Q is successful is equal to 4/5.
Whether or not these two projects are successful are independent of each other.
What is the probability that both P and Q are not successful?
-------
If they are independent, P(both not successful) = (1/3)(1/5) = 1/15
------------------------------
Cheers,
Stan H.
=====================